# Resonance RoPE: Improving Context Length Generalization of Large Language Models

### v1

@@ 7 Conclusion
We introduce Resonance RoPE, a novel enhancement of RoPE that focuses on minimizing the interpolation of RoPE features for OOD positions, thereby reducing the generalization gap and improve LLM’s performance on train-short-test-long (TSTL) scenarios. Additionally, we present a novel synthetic benchmark, PosGen, which provides a fine-grained analysis on the model’s TSTL performance regarding various token dependency patterns. Extensive experiments on our proposed PosGen and two LLM-based evaluations demonstrate Resonance RoPE’s efficacy in identifying OOD positions and its compatibility with current RoPE scaling strategies. Future work includes exploring Resonance RoPE’s performance on other foundational models, and the identification of more optimal wavelength combinations for RoPE features.
@@
Secondly, applying LLMs to long text sequences requires considerations on both performance and efficiency due to the super-linear complexity of Transformers w.r.t input length. As an improvement of the position embeddings, we focus only on improving Transformers’ performance in TSTL scenarios. An interesting future direction would be to apply Resonance RoPE to efficient Transformers for both performance and efficiency enhancements.
@@ Appendix A Proof of Theorem 1
Appendix B Detailed Experiment Settings
@@
B.1 Synthetic Task Evaluation on PosGen
For the synthetic task experiments in Section 6.1.1, we train a two-layer Transformer on each of the subtasks, with each layer following the configuration of a T5-Small model (Raffel et al., 2020). For each subtask, we train the model with different position embeddings on a training set with 10,000 sequence samples of length 64. The validation and test sets each contain 1,000 sequence samples with length 256. The sequences in the training, validation and test sets do not overlap in the first $j+k$ tokens. For all YaRN and Resonance YaRN settings, we train the model with YaRN and Resonance YaRN applied to the model with a scaling factor $s=4$, which corresponds to the TSTL setting of our evaluation. Each model is trained on each subtask for 150 epochs with a language modeling-style cross-entropy loss. Training was done with AdamW optimizer (Loshchilov and Hutter, 2019), using learning rate $2\times 10^{-4}$ and weight decay $1\times 10^{-2}$. We use a batch size of $128$ for all experiments. All hyperparameters were tuned to maximize YaRN’s validation set performance on the Semi-Recurrent subtask. All synthetic task evaluations were performed on a single NVIDIA V100 32G GPU.
B.2 LLM Evaluations
@@
  • NTK-Aware Scaling (bloc97, 2023; Xiong et al., 2023; Liu et al., 2023b), which scales the base $b$ in Equation 1 to $s\cdot b$, where $s$ is the scaling factor. We evaluate the performance without fine-tuning as used in bloc97 (2023).
  • Dynamic NTK-Aware Scaling (Peng et al., 2023; Rozière et al., 2023). This method dynamically computes the scaling factor considering the current sequence length $L_{c}$ and the original context window length $L$: $s=\frac{L_{c}}{L}$. Due to the high cost of frequently recomputing RoPE features, we evaluated its performance without fine-tuning.
  • YaRN (Peng et al., 2023). We evaluate its performance after fine-tuning.
For NTK-Aware scaling and Dynamic NTK-Aware scaling settings, we replace the original RoPE position embeddings in the model with the scaled ones and test their performance without fine-tuning following (bloc97, 2023; Peng et al., 2023). For YaRN and Resonance YaRN settings, we fine-tune the model for approximately 100M tokens on PG19’s training set (Rae et al., 2020). Our target scaled length for the 7B and 13B models is 32K and 16K, respectively, which corresponds to a scaling factor $s=8$ and $s=4$ for the position embeddings of the two models.

### v2

@@ 7 Conclusion
We introduce Resonance RoPE, a novel enhancement of RoPE that focuses on minimizing the interpolation of RoPE features for OOD positions, thereby reducing the generalization gap and improving LLM’s performance on train-short-test-long (TSTL) scenarios. Additionally, we present a novel synthetic benchmark, PosGen, which provides a fine-grained analysis of the model’s TSTL performance regarding various token dependency patterns. Extensive experiments on our proposed PosGen and two LLM-based evaluations demonstrate Resonance RoPE’s efficacy in identifying OOD positions and its compatibility with current RoPE scaling strategies. Future work includes exploring Resonance RoPE’s performance on other foundational models, and the identification of more optimal wavelength combinations for RoPE features.
@@
Secondly, applying LLMs to long text sequences requires considerations of both performance and efficiency due to the super-linear complexity of Transformers w.r.t input length. As an improvement of the position embeddings, we focus only on improving Transformers’ performance in TSTL scenarios. An interesting future direction would be to apply Resonance RoPE to efficient Transformers for both performance and efficiency enhancements.
@@ Appendix A Proof of Theorem 1
Appendix B Comparison Between Feature Gap and Embedded Vector Distance
Our proposed feature gap metric, as defined in Equation 8, shares similarities with the “embedded vector distance” metric introduced by Xiong et al. (2023):
$d(f,\hat{f})=\max_{x\in\mathcal{X}}\min_{{\begin{subarray}{c}k\in\{0,\cdots,N-1\}\\ j\in\{0,\cdots,\hat{N}-1\}\end{subarray}}}\text{dist}[f(x,k),\hat{f}(x,j)]$ (9)
where $\mathcal{X}\subset\mathbb{R}^{d}$ represents the set of vectors requiring positional embedding. This equation assesses the discrepancy in Rotary Position Embedding (RoPE) before and after a scaling operation. The distance calculation specifically compares the original RoPE, $f(\cdot,\cdot)$, to the scaled RoPE, $\hat{f}(\cdot,\cdot)$, with token positions beginning at zero. It aims to quantify the alterations in position embedding due to the scaling process.
In contrast, our feature gap metric is tailored for a more practical and common scenario, where models are trained or fine-tuned on short sequences using the already scaled RoPE embeddings. This setting emphasizes the generalization gap of the RoPE features between training and testing position ranges. The core hypothesis is that a smaller discrepancy in the RoPE features of new token positions relative to those encountered during training correlates with enhanced model generalization to novel token positions. Our metric diverges from the “embedded vector distance” in two significant aspects to better align with our use-case:
  • The distance computation shifts to compare scaled RoPE across different token positions, reflecting the operational context where training involves short sequences (train-short) and testing involves longer sequences (test-long).
  • We modify the token position ranges, $k$ and $j$, to represent token positions observed during training (in-distribution) and testing (out-of-distribution), respectively, to directly measure the generalization gap.
This adaptation of the metric allows for a more targeted evaluation of the model’s ability to generalize across different token positional distributions, which is critical in scenarios where sequence length varies significantly between training and deployment.
Appendix C Detailed Experiment Settings
@@
C.1 Synthetic Task Evaluation on PosGen
For the synthetic task experiments in Section 6.1.1, we train a two-layer Transformer on each of the subtasks, with each layer following the configuration of a T5-Small model (Raffel et al., 2020). For each subtask, we train the model with different position embeddings on a training set with 10,000 sequence samples of length 64. The validation and test sets each contains 1,000 sequence samples with length 256. The sequences in the training, validation and test sets do not overlap in the first $j+k$ tokens. For all YaRN and Resonance YaRN settings, we train the model with YaRN and Resonance YaRN applied to the model with a scaling factor $s=4$, which corresponds to the TSTL setting of our evaluation. Each model is trained on each subtask for 150 epochs with a language modeling-style cross-entropy loss. Training was done with AdamW optimizer (Loshchilov and Hutter, 2019), using learning rate $2\times 10^{-4}$ and weight decay $1\times 10^{-2}$. We use a batch size of $128$ for all experiments. All hyperparameters were tuned to maximize YaRN’s validation set performance on the Semi-Recurrent subtask. All synthetic task evaluations were performed on a single NVIDIA V100 32G GPU.
C.2 LLM Evaluations
@@
  • NTK-Aware Scaling (bloc97, 2023; Xiong et al., 2023; Liu et al., 2024), which scales the base $b$ in Equation 1 to $s\cdot b$, where $s$ is the scaling factor. We evaluate the performance without fine-tuning as used in bloc97 (2023).
  • Dynamic NTK-Aware Scaling (Peng et al., 2024; Rozière et al., 2023). This method dynamically computes the scaling factor considering the current sequence length $L_{c}$ and the original context window length $L$: $s=\frac{L_{c}}{L}$. Due to the high cost of frequently recomputing RoPE features, we evaluated its performance without fine-tuning.
  • YaRN (Peng et al., 2024). We evaluate its performance after fine-tuning.
For NTK-Aware scaling and Dynamic NTK-Aware scaling settings, we replace the original RoPE position embeddings in the model with the scaled ones and test their performance without fine-tuning following (bloc97, 2023; Peng et al., 2024). For YaRN and Resonance YaRN settings, we fine-tune the model for approximately 100M tokens on PG19’s training set (Rae et al., 2020). Our target scaled length for the 7B and 13B models is 32K and 16K, respectively, which corresponds to a scaling factor $s=8$ and $s=4$ for the position embeddings of the two models.